OpenStudy (anonymous):

Easy Math !!!!!! Art bought a 4 boxes of pens, but gave his friend Ben 8 pens, and now has a total of 112 pens. How many are in a box? Which of the following is the best equation for the problem? (Points : 1) 4X - 8 = 112 4X + 8 = 112 X /4 - 8 = 112 112 - X/4 = 8

OpenStudy (anonymous):

4x - 8 = 112 112 + 8 = 4x 120 = 4x 120 / 4 = x

OpenStudy (anonymous):

A

OpenStudy (anonymous):

We are to guide them to the correct answer, Anders, not just give them it without explanation.